• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name PU

    Log entry time 08:01:19 on October20,2002

    Entry number 88536

    keyword=Owl shift summary

    2:30 Bob Michael has worked on a Fastbus crate in the Right arm
    and fixed the problem of the deadtime.
    Now the UNBUFF mode should be used.

    Run 2566 : pedestal run

    Run 2567 : Cosmics run

    Run 2568 : Cosmics run

    We have some trouble with the acquisition. CODA was not able to discuss
    with several ROCs. However the ROCs were alive.
    First we just kill CODA and restart it. But dowload failed
    (ROCs were not responding)
    After several iterations, we decided to reset all ROCs (except TS1 and TS2)
    and restart CODA ... then it works again ! (20 min lost)

    4:45 Paul and Mark come back from downstairs. Right A1 detector now looks reasonable, but gains need to be software matched. Left gas cherenkov still looks problematic. Hopefully Bogdan can help to get this fixed during tomorrow day shift (especially if we go to machine development from 10am-noon). Left PR detector now looks good. Dead channels (except for #17 layer 1 which is OFF) in layer 1 all live now. Layer 2 low channels fixed by switching cable at Fastbus ADC to aerogel slot (slot 23 to 24) (detmap updated accordingly) or perhaps by addition of shielding (?) connector attachment at ADC. The problem went away when both of these actions were taken, so we don't know which was the fix. Software gain matching still needs to be done for PR.

    4:55 Ask for the beam

    5:00 Beam ON but MCC cannot deliver more than 20 uA. We run at 20 uA starting the spectrometer offset measurement with kinematics Ela_H01. Beam wanders all over the place, so have to wait to take data.
    6:39 Run 2580 Good run for Ela_H01 (with stable beam position) at 20 microA
    7:00 Run 2581 Optics run in Ela_H01 kinematics
    7:30 Change the kinematics from Ela_H01 to Ela_H02